# The role of porosity in the transition to inertial regime in porous media flows

### v2

@@ 4 Conclusions
The investigated physical phenomena are closely related to the tools used to quantify them. The presented results show the utility of the two discussed measures of the inertial effects in porous media – streamline-based and volume-integrated tortuosity. Each of them separately carries information about various forms of the inertial effects in the flow – $T_{s}$ is more sensitive to the changes occurring solely in the percolating volume, while in $T_{\Omega}$ the contribution of the recirculation zones is also visible. The difference between the values of the two, as well as the magnitude of the growth of $T_{\Omega}$, can be treated as an indicator of the confinement of the recirculation zones. The presented results also highlight which aspects of the inertia onset in porous media are better captured by which definition of tortuosity and why, allowing for a more sensitive choice between the two in more application-oriented problems. One may stipulate, for instance, that the information about the recirculation zones carried by the volume-integrated tortuosity might prove useful when the mixing and reaction phenomena in the porous sample are of interest. On the other hand, the exclusive contribution of the percolating volume to the streamline-based tortuosity may carry more valuable information for studies concerned with the mass transport through the sample, e.g., during gas extraction, where the gaseous phase is often transported through preferential channels and the residuals of the fluid phase occupy the dead-end pores and cavities.
5 Acknowledgments
@@
Appendix A Relation to other measures for the onset of inertia
@@
To give a better perspective on our present findings, for the discussed stochastic and simple cubic systems, we additionally calculate other measures of the inertia onset. Fig. 21 shows the relation between the mean flow velocity in the $x_{0}$-direction versus the magnitude of the forcing term (the Darcy’s law, Eq. (1)). A comparison with Figs. 11 and  16 shows that for the higher-porosity samples of each type the Reynolds number at which the mean velocity–forcing term relation becomes nonlinear corresponds roughly to the beginning of the fall of the tortuosities ($Re_{k}\approx 0.49$ and $5.75$ for the stochastic and simple cubic system, respectively). For the lower-porosity samples, the transition to the nonlinear dependence occurs later, around the minimum of $T_{\Omega}$ ($Re_{k}\approx 0.48$ and $1.45\cdot 10^{-3}$, for the stochastic and simple cubic system, respectively). Note also that for the simple cubic samples, the divergence from linearity in $\langle V_{0}\ranglegg$ relation is weaker than in the stochastic systems. This is especially in contrast to the behavior of $T_{\Omega}$ in the low-porosity SC sample, which experiences the biggest jump in value between the Darcy regime and the highest $Re_{k}$ from among all the studied systems. This suggests that the volume-integrated tortuosity might be a good indicator for the onset of inertia for extremely low-porosity systems.
@@
Figure 21: Mean flow velocity in the $x_{0}$-direction versus the magnitude of the forcing term for the stochastic systems (left) and simple cubic systems (right). The dashed lines present the first-order scaling slopes anchored at the data points with the lowest value of $g$ (for some systems, those points are not visible at the presented axes ranges). The values at the arrows show the $Re_{k}$ for the data points around which the relation becomes non-linear. The acceleration and velocity for stochastic systems are given in lattice units.
Fig 22 shows the dependence of the inertial correction $f_{c,0}$ defined as [17]
$f_{c,0}=-\left(\frac{\langle\boldsymbol{V}\rangle}{|\langle\boldsymbol{V}\rangle|\phi}-\frac{k_{0}\boldsymbol{g}}{\mu|\langle\boldsymbol{V}\rangle|}\right)\cdot\hat{\boldsymbol{x}}_{0}=-1+\frac{k_{0}g}{\mu\langle V_{0}\rangle\phi}$ (16)
@@
on the Reynolds number $Re_{k}$ for the studied stochastic and simple cubic systems. In Eq. (16), $\hat{\boldsymbol{x}}_{0}$ is the unit vector aligned with the $x_{0}$-axis, and the last equality holds under the assumption of the mean flow aligned with the $x_{0}$ direction, namely $\langle\boldsymbol{V}\rangle=\langle V_{0}\rangle\hat{\boldsymbol{x}}_{0}$. The weak inertia regime is characterized by the $f_{c,0}\propto Re_{k}^{2}$ scaling, while the strong inertia regime – by $f_{c,0}\propto Re_{k}$ scaling. In the case of stochastic samples, the latter regime is vaguely visible, which suggests that the transition between the two regimes takes place over the majority of the considered Reynolds numbers. In the case of simple cubic systems, the weak inertia regime is clearly visible for both porosities. In the several lowest $Re_{k}$ cases, the scaling becomes below-second order, which was also observed by other authors [17]. In the lower-porosity sample, the deviation from the weak inertia regime appears approximately at $Re_{k}=4\cdot 10^{-3}$, for which the $T_{\Omega}$ reaches its minimum (cf. Fig 16). For the highest Reynolds numbers, the relation changes into sublinear; however, it endures over a short range of $Re_{k}$ before the steady-state regime ends. It is possible that one observes the regime above strong inertia where $f_{c,0}\propto Re_{k}^{2}+Re_{k}$. One may relate this regime to the reduction in the increase of the slope of the volume-based tortuosity at $Re_{k}>2\cdot 10^{-2}$. For the higher-porosity simple cubic sample, the relation becomes sub-quadratic around $Re_{k}=5$, the value at which the tortuosities in Fig. 16 begin to rapidly decrease. Also in this case, for the highest Reynolds numbers, the $f_{c,0}(Re_{k})$ relation is clearly non-linear.
@@
Figure 22: The inertial corection $f_{c,0}$ (Eq. (16)) versus the Reynolds number for the stochastic systems (left) and simple cubic systems (right). The dashed lines present the first- and second-order scaling slopes for the stochastic samples and 1/21/2– and second-order scaling slopes for the simple cubic samples. The inset in the right subplot shows the magnification of the highest-$Re_{k}$ region for the lower-porosity sample.
@@ Appendix B Discretization independence study
To assess the discretization independence of $T_{\Omega}$ for the SC systems, we calculate its values from the solution of the governing equations obtained at different space discretization parameters ($h$ for uniform meshes and $h_{\text{min}}$ for refined meshes, cf. Section 2.1.3). For each porosity, the highest $Re_{k}$ is considered. In the case of the higher-porosity sample, we use the refined discretization and solve the transient problem. The results are presented in Table 1. The change of $T_{\Omega}$ between the subsequent discretizations is calculated relative to the value at the finer discretization.
@@
Table 1: Values above unity of the volume-integrated tortuosity $T_{\Omega}-1$ obtained for the highest forcing term $g$ for SC systems at various space discretization parameters used for the solution of the governing equations. The relative change in each row (rightmost column) is normalized by the value of $T_{\Omega}-1$ from the row above (one coarser discretization). The bold case row indicates the values discussed in the work.
We additionally perform the study of the independence of the $T_{s}$ value on the spacing $h_{s}$ and timestep value $\Delta t$ of the streamline generation procedure in SC samples. In this case, we take a look at the samples of the lowest $Re_{k}$, as these are the ones where the streamlines are expected to be the most tortuous, thus introducing larger errors at the segments of high curvature. We use the velocity fields obtained on the uniform $h=1/150$ or refined $h_{\text{min}}=1/150$ discretizations described in Sec. 2.1.3. We scale streamline seeding distance $h_{s}$ and the timestep for the integration along characteristics $\Delta t_{s}$ by a factor of $2$ between subsequent setups. The results are presented in Table 2
$\phi=0.1$, $Re_{k}=2.61\cdot 10^{-7}$

### v3

@@ IV Conclusions
The investigated physical phenomena are closely related to the tools used to quantify them. The presented results show the utility of the two discussed measures of the inertial effects in porous media – streamline-based and volume-integrated tortuosity. Each of them separately carries information about various forms of the inertial effects in the flow – $T_{s}$ is more sensitive to the changes occurring solely in the percolating volume, while in $T_{\Omega}$ the contribution of the recirculation zones is also visible. The difference between the values of the two, as well as the magnitude of the fall and the growth of $T_{\Omega}$, can be treated as an indicator of the confinement of the recirculation zones.
The presented results also highlight which aspects of the inertia onset in porous media are better captured by which definition of tortuosity and why, allowing for a more sensitive choice between the two in more application-oriented problems. One may stipulate, for instance, that the information about the recirculation zones carried by the volume-integrated tortuosity might prove useful when the mixing and reaction phenomena in the porous sample are of interest. On the other hand, the exclusive contribution of the percolating volume to the streamline-based tortuosity may carry more valuable information for studies concerned with the mass transport through the sample, e.g., during gas extraction, where the gaseous phase is often transported through preferential channels and the residuals of the fluid phase occupy the dead-end pores and cavities. We refer the reader to Appendix A for an additional discussion on the relation between the volume-integrated tortuosity and other, commonly used, measures of inertia in porous media flows.
V Acknowledgments
@@
VI Author contributions
D.S.: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Writing - Original Draft, Visualization
G.K.: Resources, Writing - Review & Editing, Project administration, Funding acquisition
M.M.: Conceptualization, Methodology, Resources, Writing - Review & Editing, Supervision, Project administration, Funding acquisition
VII Data availability
The velocity and pressure (density) fields generated in the present study are available at a public repository under the URL: https://doi.org/10.5281/zenodo.16794913. The repository also contains chosen scripts, source and configuration files used in the present study.
Appendix A Relation to other measures of the onset of inertia
@@
To give a better perspective on our present findings, for the discussed stochastic and simple cubic systems, we additionally calculate other measures of the inertia onset. Fig. 22 shows the relation between the mean flow velocity in the $x_{0}$-direction versus the magnitude of the forcing term (the Darcy’s law, Eq. (1)). A comparison with Figs. 7 and  15 shows that for the highest-porosity samples of each type the Reynolds number at which the mean velocity–forcing term relation becomes nonlinear corresponds roughly to the beginning of the fall of the tortuosities ($Re_{k}\approx 0.52$ and $5.84$ for the stochastic and simple cubic system, respectively). For the lowest-porosity samples, the transition to the nonlinear dependence occurs later, around the minimum of $T_{\Omega}$ ( $Re_{k}\approx 0.56$ and $4.58\cdot 10^{-3}$, for the stochastic and simple cubic system, respectively). Note also that for the simple cubic samples, the divergence from linearity in $\langle V_{0}\ranglegg$ relation is weaker than in the stochastic systems. This is especially in contrast to the behavior of $T_{\Omega}$ in the lowest-porosity SC sample, which experiences the biggest jump in value between the Darcy regime and the highest $Re_{k}$ from among all the studied systems. This suggests that the volume-integrated tortuosity might be a good indicator for the onset of inertia for extremely low-porosity systems.
@@
Figure 22: Mean flow velocity in the $x_{0}$-direction versus the magnitude of the forcing term for the stochastic systems (top) and simple cubic systems (bottom). The dashed lines present the first-order scaling slopes anchored at the data points with the lowest value of $g$ (for some systems, those points are not visible at the presented axes ranges). The values at the arrows show the accelerations around which, according to this plot, the relation becomes non-linear, corresponding to $Re_{k}=4.58\cdot 10^{-3},1.02,27.1,5.84$, in the order of the increasing porosity. The acceleration and velocity for stochastic systems are given in lattice units.
Fig 23 shows the dependence of the dimensionless inertial correction $f_{c,0}$ defined as [17]
$f_{c,0}=-\left(\frac{\langle\boldsymbol{V}\rangle}{|\langle\boldsymbol{V}\rangle|\phi}-\frac{k_{0}\boldsymbol{g}}{\nu|\langle\boldsymbol{V}\rangle|}\right)\cdot\hat{\boldsymbol{x}}_{0}=-1+\frac{k_{0}g}{\nu\langle V_{0}\rangle}$ (16)
@@
on the Reynolds number $Re_{k}$ for the studied stochastic and simple cubic systems. In Eq. (16), $\hat{\boldsymbol{x}}_{0}$ is the unit vector aligned with the $x_{0}$-axis, and the last equality holds under the assumption of the mean flow aligned with the $x_{0}$ direction, namely $\langle\boldsymbol{V}\rangle=\langle V_{0}\rangle\hat{\boldsymbol{x}}_{0}$. The weak inertia regime is characterized by the $f_{c,0}\propto Re_{k}^{2}$ scaling, while the strong inertia regime – by $f_{c,0}\propto Re_{k}$ scaling. In the case of stochastic samples, the latter regime is vaguely visible, which suggests that the transition between the two regimes takes place over the majority of the considered Reynolds numbers. In the case of simple cubic systems, the weak inertia regime is clearly visible for both porosities. In the several lowest $Re_{k}$ cases for each porosity, the scaling becomes below-second order, which was also observed by other authors [17]. In the lowest-porosity sample, the deviation from the weak inertia regime appears approximately at $Re_{k}=10^{-2}$, for which the $T_{\Omega}$ reaches its minimum (cf. Fig 15). For the highest Reynolds numbers, the relation changes into sublinear; however, it endures over a short range of $Re_{k}$ before the steady-state regime ends. It is possible that one observes the regime above strong inertia where $f_{c,0}\propto Re_{k}^{2}+Re_{k}$. One may relate this regime to the reduction of the slope of the volume-based tortuosity at $Re_{k}>6\cdot 10^{-2}$. As the porosity increases, the deviation from the quadratic scaling of $f_{c,0}$ shifts towards the lower Reynolds numbers, closer to the start of the $T_{\Omega}$ decrease – $Re_{k}=0.3,\>2,\>5$ for porosities $\phi=0.59,\>0.93,\>0.999$, respectively. This, along with the similar observations made for the deviation from Darcy’s law, supports the idea that in low-porosity samples, the onset of inertia manifests itself rather through the reorganization of the flow paths (e.g., straightening of the streamlines) than significant changes in the averaged transport coefficients (like the apparent permeability or the magnitude of the inertial correction.) Also, in all SC samples, for the highest Reynolds numbers, the $f_{c,0}(Re_{k})$ relation is clearly non-linear. We note that if one considered the dimensional inertial correction, $f_{i,0}\propto\langle V_{0}\rangle f_{c,0}$, the second order scaling $f_{c,0}\propto Re_{k}^{2}$ would turn into the cubic scaling $f_{i,0}\propto Re_{k}^{3}$ which would directly highlight the dominance of the $\langle V_{0}\rangle^{3}$ term in the corrections to the Darcy’s law, Eq. (2).
@@
Figure 23: The inertial correction $f_{c,0}$ (Eq. (16)) versus the Reynolds number for the stochastic systems (top) and simple cubic systems (bottom). The dashed lines present the first- and second-order scaling slopes for the stochastic samples and 1/21/2– and second-order scaling slopes for the simple cubic samples.
Finally, to allow for easy comparison with other works, in Table 3 we provide the conversion factors from $Re_{k}$ used in the present study, to the Reynolds number $Re_{d}$, which uses the grain size (diameter in case of spheres and side length in case of cubes) as the reference length. In Fig. 24 we show the same data as in the bottom subplot of Fig. 14 but plotted against $Re_{d}$.
System type $\phi$ $L_{\text{ref}}$ $Re_{d}/Re_{k}$
Stochastic $0.7$ 1919 $5.99$
$0.9$ $10$ $1.82$
SC $0.1$ $1.31$ $300$
$0.59$ $0.92$ $12.8$
$0.93$ $0.52$ $2.01$
$0.999$ $0.124$ $0.148$
Table 3: Conversion factors from the Reynolds number used in the present study, $Re_{k}$, to the Reynolds number $Re_{d}$ using the size of the grain as the reference length.
Figure 24: Volume-integrated tortuosity for all SC samples rescaled by $T_{\Omega}$ at vanishingly small Reynolds number. Note that the horizontal axis shows $Re_{d}$, not $Re_{k}$.
@@ Appendix B Discretization independence study
To assess the discretization independence of $T_{\Omega}$ for the SC systems, we calculate its values from the solution of the governing equations obtained at different space discretization parameters ($h$ for uniform meshes and $h_{\text{min}}$ for refined meshes, cf. Section II.1.3). For each porosity, the highest $Re_{k}$ is considered. In the case of the highest-porosity sample, we use the refined discretization and solve the transient problem. The results are presented in Table 4. The change of $T_{\Omega}$ between the subsequent discretizations is calculated relative to the value at the finer discretization. The results show that increasing the discretization density by a factor of $1.3$ from the one used to produce the data presented in the Results section is not expected to alter the value of $T_{\Omega}-1$ by more than $3\%$.
@@
$\phi=0.59$, $g=3\cdot 10^{4}$
$h$ $T_{\Omega}-1$ $\%$ change
$\mathbf{1/150}^{*}$ $\mathbf{2.832\cdot 10^{-2}}$
$1/195$ $2.848\cdot 10^{-2}$ $0.565$
$1/253$ $2.853\cdot 10^{-2}$ $0.193$
$\phi=0.93$, $g=5.6\cdot 10^{3}$
$h$ $T_{\Omega}-1$ $\%$ change
$1/115$ $8.451\cdot 10^{-3}$
$\mathbf{1/150}$ $\mathbf{8.432\cdot 10^{-3}}$ 0.232
$1/195$ $8.398\cdot 10^{-3}$ $0.405$
@@
Table 4: Values above unity of the volume-integrated tortuosity, $T_{\Omega}-1$, obtained for the highest forcing term $g$ for SC systems at various space discretization parameters used for the solution of the governing equations. The relative change in each row (rightmost column) is calculated with respect to the value of $T_{\Omega}-1$ from the row above (one coarser discretization). The bold case row indicates the values discussed in the work. The asterisk denotes the data obtained from pimpleFoam simulation with $\Delta t=3.5\cdot 10^{-6}$.
We additionally perform the study of the independence of the $T_{s}$ value on the spacing $h_{s}$ and timestep value $\Delta t_{s}$ of the streamline generation procedure in SC samples. In this case, we take a look at the samples of the lowest $Re_{k}$, as these are the ones where the streamlines are expected to be the most tortuous, thus introducing larger errors at the segments of high curvature. We use the velocity fields obtained on the uniform $h=1/150$ or refined $h_{\text{min}}=1/150$ discretizations described in Sec. II.1.3. We scale streamline seeding distance $h_{s}$ and the timestep for the integration along characteristics $\Delta t_{s}$ by a factor of $2$ between subsequent setups. The results are presented in Table 5. Also here, doubling the discretizations density is not expected to change $T_{s}-1$ by more than $3\%$.
$\phi=0.1$, $Re_{k}=8.24\cdot 10^{-7}$
@@
$\phi=0.59$, $Re_{k}=3.72\cdot 10^{-14}$
$h_{s}$ $\Delta t_{s}$ $T_{s}-1$ $\%$ change
$\cdot 10^{-2} 11$$\cdot 10^{-2}$ $1.766\cdot 10^{-2}$
$5\cdot 10^{-3}$ $5\cdot 10^{-3}$ $1.792\cdot 10^{-2}$ $1.454$
$\mathbf{2.5\cdot 10^{-3}}$ $\mathbf{2.5\cdot 10^{-3}}$ $\mathbf{1.805\cdot 10^{-2}}$ $\mathbf{0.724}$
$1.25\cdot 10^{-3}$ $1.25\cdot 10^{-3}$ $1.819\cdot 10^{-2}$ $0.785$
$\phi=0.93$, $Re_{k}=1.73\cdot 10^{-12}$
$h_{s}$ $\Delta t_{s}$ $T_{s}-1$ $\%$ change
$10^{-2}$ $10^{-2}$ $8.076\cdot 10^{-3}$
$\mathbf{5\cdot 10^{-3}}$ $\mathbf{5\cdot 10^{-3}}$ $\mathbf{8.156\cdot 10^{-3}}$ $\mathbf{0.992}$
$2.5\cdot 10^{-3}$ $2.5\cdot 10^{-3}$ $8.196\cdot 10^{-3}$ $0.495$
$1.25\cdot 10^{-3}$ $1.25\cdot 10^{-3}$ $8.240\cdot 10^{-3}$ $0.530$